OpenStudy (anonymous):

Which of the following ordered pairs is not a solution to the inequality y -3x - 2? A:(1, 1) B:(0, 0) C:(-1, 0) D:(0, -1)

OpenStudy (anonymous):

you say inequality, but I don't see any sign like.. > < \[\leq \;\;\; \geq\]

OpenStudy (anonymous):

o my bad theres a \[\ge\] in between y and -3x-2